OpenStudy (anonymous):

How do you factor quadratics using grouping? An example would be :4x^2-9x+5=0 I'm having difficulty understand this concept and would appreciate your help

hartnn (hartnn):

@ww2greatbritainuk Hi, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (amistre64):

consider 2 factors: (x+a)(x+b) when we multiply them together ... use your favorite method. we get: x^2 + ax + bx + ab , which simplifies to x^2 + (a+b)x + ab

OpenStudy (amistre64):

the idea with grouping is to try to split the middle (a+b) term into a sum of the factors of the last term (ab)

OpenStudy (anonymous):

So 5/4 and 1?

OpenStudy (anonymous):

(4x^2-4x)(-5x+5) 4x(x-1)-5(x-1)=0 (4x-5)(x-1)=0 x=5/4 x= 1

OpenStudy (amistre64):

i was never great at grouping an usually go a different route to factor :4x^2-9x+5=0 4(5) = 20, last term is + so factors are same sign middle term is -, so factors are both negative 20 = 5*4, and 5+4 = 9 therefore (x-5/4) (x-4/4) fit my idea, reduce and move denominators back in front (4x-5) (x-1)

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

youre welcome :)